# A Scalable Diagonalization Framework for Tensor-Product Bitstring Selected Configuration Interaction

### 05 Mar 2026 v4

@@ III.2 Compactness of the TPB representation from SCI-derived bitstrings
The same bitstring-selection procedure was further applied to CN, Cr2, and N2 (cc-pVTZ), with the results summarized in Table 4. For CN, several geometries along the PES were examined, while single-point calculations were performed for Cr2 and N2 (cc-pVTZ). Across all systems, TBSCI systematically approaches the FCI limit as the bitstring selection threshold $\delta$ is reduced. For the molecular systems considered here, sub-millihartree accuracy is achieved at $\delta=10^{-9}$ while using well below $0.56\%$ of the FCI determinants. Although the required fraction $N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}$ is larger for the strongly correlated transition-metal system Cr2 than for main-group molecules, near-FCI accuracy is still recovered with a substantially reduced determinant space. Reference FCI energies for CN with cc-pVDZ, cc-pVTZ, and complete basis set (CBS) extrapolation along the PES are listed in Table S3 of the Supplementary Material.
@@
Cr2     $R=1.5$ Å    $E_{\mathrm{HF}}=-2086.572971$     $E_{\mathrm{DMRG}}=-2086.44478$
$1\times 10^{-6}$ 5,669 5,994,031 41.229 (12, 11.2)
$1\times 10^{-7}$ 20,756 71,784,176 19.038 (12, 180.9)
$1\times 10^{-8}$ 65,895 679,551,775 8.368 (12, 2409.7)
$1\times 10^{-9}$ 181,976 5,061,656,122 3.752 (48, 6770.8)
$1\times 10^{-10}$ 427,692 27,568,299,482 1.867 (386, 7228.9)
0 1,722,583 490,574,659,035 0.915 (7200, 10395.3)
N2     $R=2.068$ a.u.     $E_{\mathrm{HF}}=-108.991735$     $E_{\mathrm{FCIQMC}}=-109.40251$
@@ IV Conclusions and Prospects
Looking forward, the next stage of development concerns refining determinant selection within the TPB representation to further compress the resulting TBSCI wavefunction. One possible route is to apply FCIQMC-style stochastic sampling [29, 28] within the TPB determinant space, using walker populations as a practical proxy for determinant importance. A second direction involves iteration-adaptive redistribution of CI-vector segments based on measured wall times from preceding Davidson iterations, with the aim of reducing residual load imbalance. Finally, extending the distributed framework to incorporate second-order perturbative corrections will be essential for achieving higher accuracy in large-scale applications.

### 29 Apr 2026 v5

@@ III.2 Compactness of the TPB representation from SCI-derived bitstrings
The same bitstring-selection procedure was further applied to CN, Cr2, and N2 (cc-pVTZ), with the results summarized in Table 4. For CN, several geometries along the PES were examined, while single-point calculations were performed for Cr2 and N2 (cc-pVTZ). Across all systems, TBSCI systematically approaches the FCI limit as the bitstring selection threshold $\delta$ is reduced. For the molecular systems considered here, sub-millihartree accuracy is achieved at $\delta=10^{-10}$ while using well below $0.56\%$ of the FCI determinants. Although the required fraction $N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}$ is larger for the strongly correlated transition-metal system Cr2 than for main-group molecules, near-FCI accuracy is still recovered with a substantially reduced determinant space. Reference FCI energies for CN with cc-pVDZ, cc-pVTZ, and complete basis set (CBS) extrapolation along the PES are listed in Table S3 of the Supplementary Material.
@@
Cr2     $R=1.5$ Å    $E_{\mathrm{HF}}=-2086.572971$     $E_{\mathrm{DMRG}}=-2086.444784$
$1\times 10^{-6}$ 5,669 5,994,031 41.233 (12, 11.2)
$1\times 10^{-7}$ 20,756 71,784,176 19.042 (12, 180.9)
$1\times 10^{-8}$ 65,895 679,551,775 8.372 (12, 2409.7)
$1\times 10^{-9}$ 181,976 5,061,656,122 3.756 (48, 6770.8)
$1\times 10^{-10}$ 427,692 27,568,299,482 1.871 (386, 7228.9)
0 1,722,583 490,574,659,035 0.919 (7200, 10279.3)
N2     $R=2.068$ a.u.     $E_{\mathrm{HF}}=-108.991735$     $E_{\mathrm{FCIQMC}}=-109.40521$
@@ IV Conclusions and Prospects
Looking forward, the next stage of development concerns refining determinant selection within the TPB representation to further compress the resulting TBSCI wavefunction. One possible route is to apply FCIQMC-style stochastic sampling [36, 35] within the TPB determinant space, using walker populations as a practical proxy for determinant importance. In the present work, all calculations were performed using canonical Hartree–Fock orbitals. Since natural orbitals [37, 38] and split-localized orbitals [39] have been shown to provide more compact CI expansions, incorporating such optimized orbitals into the TPB representation is expected to further enhance its compactness and will be explored in future work. Another important direction concerns the incorporation of second-order perturbative corrections (PT2). In principle, PT2 may be formulated on top of the TBSCI wavefunction. As a non-iterative procedure, PT2 does not require explicit storage of the full interacting determinant space, and the contributing determinants can be generated and evaluated in batches. Achieving this while maintaining high parallel efficiency and mitigating communication overhead will require further algorithmic developments and program optimization, potentially including semi-stochastic strategies. [14, 15] Finally, further improvements in parallel efficiency will be pursued, for example through mixed-precision implementations to reduce communication volume and iteration-adaptive redistribution of CI-vector segments based on measured wall times from preceding Davidson iterations to reduce residual load imbalance.